OpenStudy (anonymous):

(3w-27)/(14) ÷ (w-9)/(49w) Divide and simplify.

OpenStudy (anonymous):

\[\frac{3(w-9)}{14}\times\frac{49w}{w-9}\]w-9 cancle\[\frac{3\times49w}{14}\]\[\frac{147w}{14}\]\[\frac{14w}{2}\]

OpenStudy (anonymous):

sorry 21w\2

OpenStudy (anonymous):

i was wrong it is 3 ( w - 9 ) x 49w ------------ ------ 14 w - 9 the w-9 s cancel out and 49w becomes 7w and 14 becomes 2 as they are both divisible by 7 so that gives 147 w -------- 2

OpenStudy (anonymous):

sorry 21w / 2 typo

OpenStudy (anonymous):

\[\frac{21w}{2}\]

OpenStudy (anonymous):

we al keep tye wrong sorry